OpenStudy (anonymous):

Please check my work! (easy rate of change question) The function below shows the number of car owners f(t), in thousands, in a city in different years t: f(t) = 0.25t2 − 1.5t + 2.5 The average rate of change of f(t) from t = 5 to t = 8 is ______ thousand owners per year.

OpenStudy (anonymous):

I got 1.75

OpenStudy (anonymous):

First, I plugged 5 into the equation and I got 1.25 Then I plugged in 8 and got 6.5 Then I used the average rate of change formula and got: \[\frac{ 1.25 - 6.5 }{ 5 - 8 } = \frac{ -5.25 }{ -3 } = 1.75\]

OpenStudy (welshfella):

yes that's correct

OpenStudy (anonymous):

Thanks!

OpenStudy (welshfella):

yw